# Danoliteracy of Generative, Large Language Models

### v1

@@ 6 Concerns of the Ethical Impacts
The results presented here only focus on model capabilities but, on the leaderboard site, versions of other important dimensions for model applicability are presented; such as model efficiency, model likelihood calibration and model generated output toxicity. However, these are presented with a disclaimer as preliminary results and our work on other crucial dimensions such as GLLM performance fairness across gender and nationalities or robustness to input noise have not been released due to limitations to current datasets to robustly carry out these analyses.
There is an increased risk of bias, fairness and toxicity violations in low-resource languages to which models are less tuned. Problematically, when the evaluation situation is also low-resource, these risks might be undiscovered for practitioners that only focus on a model capabilities. Further work is crucially needed but for now, the leaderboard site displays a disclaimer against blindly trusting that high benchmark numbers mean predictable downstream performance or applying GLLMs with unchecked assumptions about robustness, fairness, bias, and toxicity.
@@ 7 Limitations
The presented benchmark consists of eight specific scenarios: Although we find high scenario result correlation, all our statements about model performance on Danish in general are evidently biased by the scenario selection. A similar statement can be made about prompt and metric design decisions though these seem robust in ablation studies in Appendix A.
@@ Acknowledgments
This work was supported by the Pioneer Centre for AI, DNRF grant number P1.
@@ 73. Tyrkiet
Table 2: Alternative prompting and scoring approaches to the Citizenship Test run for a subset of models including a baseline outputting a fixed, random string and Danish-tuned versions of base GLLMs. Std. is the prompt version presented in the benchmark, Figure 11, Simple Q removes the first instruction and the markdown headers, simply presenting the question, the options and the final text. No options asks the question openly without multiple-choice options, choosing argmax similarity score, as in Section 3.1, as model choice. Presented with 95% Wald confidence interval.
Std. Detailed Instructions
@@
Danoliterate Mistral 7B $57\pm 3$ $59\pm 3$
Danoliterate LlaMa 2 7B $52\pm 3$ $58\pm 3$
@@ D.2 Factor Analysis
For the EFA on the $54\times 8$ scenario results, the Bartlett Sphericity Bartlett (1951) $p$ value is $<2\cdot 10^{-16}$ and the Kaiser-Meyer-Olkin Test Kaiser (1970) yields a variance proportion of $90\%$, both suggesting that the data is usable for EFA. Fitting an EFA using the Scikit-learn Factor Analysis model yields $\lambda_{1}=5.9,\lambda_{2}=0.3$. Explained factor variance is calculated as eigenvalue proportion of summed eigenvalues, and the analysis is repeated for scenario results acquired from the open API at scandeval.com/danish-nlg/, at crfm.stanford.edu/helm/lite/v1.3.0/#/leaderboard, and using the OpenLLM Leaderboard Scraper GitHub project999github.com/Weyaxi/scrape-open-llm-leaderboard. The datasets updated to most recent versions on June 12th, 2024.

### v2

@@ 6 Concerns of the Ethical Impacts
The results presented here only focus on model capability but, on the leaderboard site, versions of other important dimensions for model applicability are presented; such as model efficiency, model likelihood calibration and model generated output toxicity. However, these are presented with a disclaimer as preliminary results and our work on other crucial dimensions such as GLLM performance fairness across gender and nationalities or robustness to input noise have not been released due to limitations to current datasets to robustly carry out these analyses.
There is an increased risk of bias, fairness and toxicity violations in low-resource languages to which models are less tuned. Problematically, when the evaluation situation is also low-resource, these risks might be undiscovered for practitioners that only focus on a model capability. Further work is crucially needed but for now, the leaderboard site displays a disclaimer against blindly trusting that high benchmark numbers mean predictable downstream performance or applying GLLMs with unchecked assumptions about robustness, fairness, bias, and toxicity.
@@ 7 Limitations
The presented benchmark consists of eight specific scenarios: Although we find high correlation between scenario results, all our statements about model performance on Danish in general are evidently biased by the scenario selection. A similar statement can be made about prompt and metric design decisions though these seem robust in ablation studies in Appendix A.
@@ Acknowledgments
We would like to thank all the anonymous reviewers for the insightful and helpful comments. This work was supported by the Pioneer Centre for AI, DNRF grant number P1.
@@ 73. Tyrkiet
Table 2: Alternative prompting and scoring approaches to the Citizenship Test run for a subset of models including a baseline outputting a fixed, random string and Danish-tuned versions of base GLLMs. Std. is the prompt version presented in the benchmark, Figure 11, Simple Q removes the first instruction and the markdown headers, simply presenting the question, the options and the final text. No opt. asks the question openly without multiple-choice options, choosing argmax similarity score, as in Section 3.2, as model choice. Presented with 95% Wald confidence interval.
Std. Detailed
@@
Dano. Mistral 7B $57\pm 3$ $59\pm 3$
Dano. LlaMa 2 7B $52\pm 3$ $58\pm 3$
@@ D.2 Factor Analysis
For the EFA on the $83\times 8$ scenario results, the Bartlett Sphericity Bartlett (1951) $p$ value is $<2\cdot 10^{-16}$ and the Kaiser-Meyer-Olkin Test Kaiser (1970) yields a variance proportion of $90\%$, both suggesting that the data is usable for EFA. Fitting an EFA using the Scikit-learn Factor Analysis model yields $\lambda_{1}=5.9,\lambda_{2}=0.3$. Explained factor variance is calculated as eigenvalue proportion of summed eigenvalues, and the analysis is repeated for scenario results acquired from the open API at scandeval.com/danish-nlg/, at crfm.stanford.edu/helm/lite/v1.3.0/#/leaderboard, and using the OpenLLM Leaderboard Scraper GitHub project101010github.com/Weyaxi/scrape-open-llm-leaderboard. The datasets updated to most recent versions on January 13th, 2025.